OpenStudy (anonymous):

The point (10,1) lies on a circle and the center of the circle is at the point (7,1). What is the equation of this circle?

OpenStudy (anonymous):

What is the radius? You can find it by drawing a line from (10,1) to (7,1) and finding the length of that line. Then, a circle is defined as x^2 + y^2 = r^2 where r = radius.